Find class boundaries , midpoint, and width for the class. 13.3-16.7
1
Expert's answer
2020-09-13T17:37:58-0400

Lower Boundary: 13.3-0.05=13.25.

Upper Boundary: 16.7+0.05=16.75.


Midpoint: "\frac{13.3+16.7}{2}=15."


Width: "16.75-13.25=3.5."